an open link chain is there and each link is of mass .1 kg........it is being pulled against gravity in vertically upward direction....find the force on the link1(the lowest one)by link 2...by 3 on 2......by 4 on 3.......by 4 on 5..........

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

Let the force F is aplied on the link and it is moving with and acceleration a.

and  chain has total 1 link

F- n*1*g =na

a= F/n -g

 

let N1 force is applied by link 2 on link 1

 N1- 1*g = 1*a

N1 = g+a

  =F/n

let N2 force is applied by link 3 an link 2

N2 -2*g =2*a

 N2 = 2(g+a)

        =2F/n

similerly other link

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin